### v2

# Making Non-Negative Polynomials into Sums of Squares

Abstract
We investigate linear operators $A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$. We give explicit operators $A$ such that, for fixed $d\in\mathds{N}_{0}$ and closed $K\subseteq\mathds{R}^{n}$, $e^{A}\mathrm{Pos}(K)_{\leq 2d}\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$. We give an explicit operator $A$ such that $e^{A}\mathrm{Pos}(\mathds{R}^{n})\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. For $K\subseteq\mathds{R}^{n}$, we give a condition such that $A$ exists with $e^{A}\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. For $K=[0,\infty)^{n}$, this condition is fulfilled. We show that, for compact $K\subseteq\mathds{R}^{n}$, there is no bijective linear operator $T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ with $T\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. In the framework of regular Fréchet Lie groups and Lie algebras we investigate the linear operators $A$ such that $e^{tA}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ is well-defined for all $t\in\mathds{R}$. We give a three-line-proof of Stochel’s Theorem.
keywords:
non-negative polynomials, sums of squares, linear operator, regular Fréchet Lie group, Positivstellensatz, semi-definite optimization
MSC:
[2010] Primary 11E25; Secondary 13J30, 44A60, 47A57, 90C22.
journal: ArXiv
1 Introduction
Let $n\in\mathds{N}$, $K\subseteq\mathds{R}^{n}$ be closed, and $\mathds{R}[x_{1},\dots,x_{n}]$ be the polynomials in $n$ variables with real coefficients. The set
$\mathrm{Pos}(K):=\big{\{}f\in\mathds{R}[x_{1},\dots,x_{n}]\,\big{|}\,f\geq 0\ \text{on}\ K\big{\}}$
of all polynomials which are non-negative on $K$ belongs to the most important structures in mathematics and applications [1, 15, 13, 26].
Let $\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$ be the cone of sums of squares. For a compact semi-algebraic set $K\subseteq\mathds{R}^{n}$ defined by finitely many polynomial inequalities $g_{i}(x)\geq 0$,
$K:=\{x\in\mathds{R}^{n}\,|\,g_{1}(x)\geq 0,\ \dots,\ g_{m}(x)\geq 0\},$
a polynomial $p$ which is strictly positive on $K$, i.e., $p(x)>0$ for all $x\in K$, has the representation
$p=\sum_{e=(e_{1},\dots,e_{m})\in\{0,1\}^{m}}g_{1}^{e_{1}}\cdots g_{m}^{e_{m}}\cdot\sigma_{e}$ (1)
with $\sigma_{e}\in\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. This was proved by Konrad Schmüdgen [23]. Another representation is
$p=\sigma_{0}+\sum_{e=1}^{m}g_{e}\cdot\sigma_{e}$ (2)
with $\sigma_{e}\in\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$ which was proved by Mihai Putinar [22]. See [1, 15, 26] for more details.
Both representations (1) and (2) only hold for $p>0$ on $K$. If
$\min_{x\in K}p_{i}(x)\quad\xrightarrow{i\to\infty}\quad 0,$
then
$\max_{e}\deg\sigma_{i,e}\quad\xrightarrow{i\to\infty}\quad\infty$
even when $p_{i}\to p$ and $\deg p_{i},\deg p\leq d<\infty$. This degree blow up causes serious troubles in optimization, since even when an optimum is attained at a polynomial $p$ of finite degree, the implementation using (1) or (2) blows up. Consequently, even in the simplest case of calculating
$\min_{p\in S}L(p)$ (3)
with a linear functional $L:\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}\to\mathds{R}$ and $S\subseteq\mathrm{Pos}(K)_{\leq d}$ the semi-definite program with (1) or (2) must in general exceed the finite degree $d$ of the problem (3). Several works investigated these degree blow ups, see e.g. [21, 24, 18, 10, 14, 4, 6] and references therein. Hence, it is clear that the representations (1) and (2) are broken beyond repair. And what is broken and can not be fixed must be replaced.
In the current work we investigate linear operators
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
especially in
$e^{tA}:=\sum_{k\in\mathds{N}_{0}}\frac{t^{k}\cdot A^{k}}{k!}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
with $t\in\mathds{R}$. The operators $A$ are investigated to determine if a given set $S\subseteq\mathrm{Pos}(K)$ can be moved by $e^{A}$ to a nicer set $e^{A}S$, e.g. in the sums of squares. This results in a transformation of the problem (3) to
$\min_{p\in S}L(p)=\min_{p\in e^{A}S}\tilde{L}(p)$
with the linear functional $\tilde{L}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}$ defined by $\tilde{L}(p):=L(e^{-A}p)$.
Structure of the Paper
In the following Section 2 we give short preliminaries of Fréchet spaces, LF-spaces, and regular Fréchet Lie groups to make the work as self-contained as possible. In Section 3 we present our main results (Theorems 3.2, 3.5, and 3.11, and 3.8), how to make non-negative polynomials into sums of squares, and how this is applied to semi-definite optimization (3.3). In 3.11 we show that, for compact $K\subseteq\mathds{R}^{n}$, there is no bijective linear operator $T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ with $T\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. In Section 4 we investigate the set $\mathfrak{g}$ of all linear maps $A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ such that $e^{tA}$ is well-defined, especially the structure of regular Fréchet Lie algebras in $\mathfrak{g}$. In Section 5 we give a three-line-proof of Stochel’s Theorem [27]. In Section 6 we give an open problem.
2 Preliminaries
For a set $S$ we denote by $\mathrm{int}\,S$ its interior.
A Fréchet space is a metrizable, complete, and locally convex topological vector space [28]. Our leading example is the set of real sequences $\mathds{R}^{\mathds{N}_{0}^{n}}$. Here, the topology is generated by the semi-norms $\|\cdot\|_{d}$ defined by
$\|s\|_{d}:=\max_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|\leq d}|s_{\alpha}|$
for every $d\in\mathds{N}_{0}$ and sequence $s=(s_{\alpha})_{\alpha\in\mathds{N}_{0}^{n}}$ with $|\alpha|:=\alpha_{1}+\dots+\alpha_{n}$ for all $\alpha=(\alpha_{1},\dots,\alpha_{n})\in\mathds{N}_{0}^{n}$. To understand the Fréchet topology of $\mathds{R}^{\mathds{N}_{0}^{n}}$ it is sufficient to understand the convergence in this space. Let $s^{(k)}=(s_{k,\alpha})_{\alpha\in\mathds{N}_{0}^{n}}$ for all $k\in\mathds{N}$. Then
$s^{(k)}\xrightarrow{k\to\infty}s\qquad\text{iff}\qquad s_{k,\alpha}\xrightarrow{k\to\infty}s_{\alpha}\ \text{for all}\ \alpha\in\mathds{N}_{0}^{n},$
i.e., $\mathds{R}^{\mathds{N}_{0}^{n}}$ has the coordinate-wise convergence.
The space of polynomials $\mathds{R}[x_{1},\dots,x_{n}]$ is a LF-space and its topology can be described by the convergence of $(p_{k})_{k\in\mathds{N}}\subseteq\mathds{R}[x_{1},\dots,x_{n}]$:
$p_{k}\xrightarrow{k\to\infty}p\qquad\text{iff}\qquad\sup_{k\in\mathds{N}}\deg p_{k}=D<\infty\ \text{and}\ p_{k}\to p\ \text{in}\ \mathds{R}[x_{1},\dots,x_{n}]_{\leq D}.$
The space $\mathds{R}^{\mathds{N}_{0}^{n}}$ and $\mathds{R}[x_{1},\dots,x_{n}]$ are dual to each other with the pairing
$\langle s,p\rangle:=L_{s}(p)$
where $s\in\mathds{R}^{\mathds{N}_{0}^{n}}$, $p\in\mathds{R}[x_{1},\dots,x_{n}]$, and $L_{s}$ is the Riesz functional
$L_{s}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},\qquad L_{s}(x^{\alpha}):=s_{\alpha}.$
All linear operators $T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ have a representation
$T=\sum_{\alpha\in\mathds{N}_{0}^{n}}q_{\alpha}\cdot\partial^{\alpha}$
with unique coefficients $q_{\alpha}\in\mathds{R}[x_{1},\dots,x_{n}]$. A sequence $(T_{k})_{k\in\mathds{N}}$ of such linear operators converges to a linear operator $T$ iff $q_{k,\alpha}\to q_{\alpha}$ for all $\alpha\in\mathds{N}_{0}^{n}$ or equivalently iff $T_{k}p\to Tp$ for all $p\in\mathds{R}[x_{1},\dots,x_{n}]$.
Since $\mathds{R}[x_{1},\dots,x_{n}]$ is a nuclear space, by the Schwartz Kernel Theorem [28, Ch. 51], every linear operator $T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ also has the representation
$T=\sum_{i\in\mathds{N}_{0}}l_{i}\cdot f_{i}$
with linear functionals $l_{i}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}$ and $f_{i}\in\mathds{R}[x_{1},\dots,x_{n}]$ for all $i\in\mathds{N}_{0}$.
We assume the reader has a basic knowledge about Lie groups and Lie algebras [29]. We only give the definition of their infinite dimensional version.
Definition 2.1 (see e.g. [20, p. 63, Dfn. 1.1]).
We call $(G,\,\cdot\,)$ a regular Fréchet Lie group if the following conditions are fulfilled:
  1. (i)
    $G$ is an infinite dimensional smooth Fréchet manifold.
  2. (ii)
    $(G,\,\cdot\,)$ is a group.
  3. (iii)
    The map $G\times G\to G$, $(A,B)\mapsto A\cdot B^{-1}$ is smooth.
  4. (iv)
    The Fréchet Lie algebra $\mathfrak{g}$ of $G$ is isomorphic to the tangent space $T_{e}G$ of $G$ at the unit element $e\in G$.
  5. (v)
    $\exp:\mathfrak{g}\to G$ is a smooth mapping such that
    $\frac{\mathrm{d}}{\mathrm{d}t}\exp(tu)\big{|}_{t=0}=u$
    for all $u\in\mathfrak{g}$.
  6. (vi)
    The space $C^{1}(G,\mathfrak{g})$ of $C^{1}$-curves in $G$ coincides with the set of all $C^{1}$-curves in $G$ under the Fréchet topology.
For more on infinite dimensional calculus, regular Fréchet Lie groups, and their Lie algebras see e.g. [19, 20, 25].
3 Making Non-Negative Polynomials into Sums of Squares
The following is our first main result.
Theorem 3.2.
Let $n\in\mathds{N}$, $d\in\mathds{N}_{0}$, $S\subseteq\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}$, and let $C\subseteq\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}$ be a full dimensional cone such that there is a linear functional
$l:\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}\to\mathds{R}\qquad\text{with}\qquad l(p)>0\quad\text{for all}\quad p\in(S\cup C)\setminus\{0\}.$
If $f\in\mathrm{int}\,C$ and the linear operator $A$ is defined by
$A:\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq d},\quad p\mapsto Ap:=l(p)\cdot f,$
then there exists a constant $\tau=\tau(f,C,S,l)>0$ such that
$e^{tA}S\quad\subseteq\quad C$
for all $t\geq\tau$.
Proof.
Let
$\tilde{S}:=\overline{\mathrm{cone}\,\mathrm{conv}\,S}$
be the closed convex cone generated by $S$. Since $l(p)>0$ for all $p\in S$, $\tilde{S}$ is closed and pointed, i.e., $\tilde{S}$ has a compact base $B$.
Let $p\in B$ and $f\in\mathrm{int}\,C$, i.e., $l(p)>0$ and $l(f)>0$. Then
$Ap=l(p)\cdot f\qquad\text{and}\qquad A^{k}p=l(p)\cdot l(f)^{k-1}\cdot f$
for all $k\in\mathds{N}$. Hence,
$e^{tA}p=\sum_{k\in\mathds{N}_{0}}\frac{t^{k}}{k!}\cdot A^{k}p=p+l(p)\cdot\sum_{k\in\mathds{N}}\frac{t^{k}}{k!}\cdot l(f)^{k-1}\cdot f=p+\frac{l(p)}{l(f)}\cdot\left(e^{t\cdot l(f)}-1\right)\cdot f$
and
$\frac{l(f)}{l(p)}\cdot\left(e^{t\cdot l(f)}-1\right)^{-1}\cdot e^{tA}p\quad\xrightarrow{t\to\infty}\quad f.$
Since
$\frac{l(f)}{l(p)}\cdot\left(e^{t\cdot l(f)}-1\right)^{-1}\cdot e^{tA}p$
is continuous in $t$ and $p$ and since $f$ is an inner point of $C$, there exists a neighborhood $U\subseteq C$ of $f$ and a constant $\tau\geq 0$ such that
$e^{tA}S\quad\subseteq\quad e^{tA}\tilde{S}\quad\subseteq\quad\mathrm{cone}\,U\quad\subseteq\quad C$
for all $t\geq\tau$. ∎
In the previous theorem it is clear that the main space $\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}$ can be replaced by any Banach space $\mathcal{V}$. For special and important cases of $S$ and $C$ in 3.2 we get the following corollary.
Corollary 3.3.
Let $n\in\mathds{N}$, $d\in\mathds{N}_{0}$, $K\subseteq\mathds{R}^{n}$ be closed, and let $S\subseteq\mathrm{Pos}(K)_{\leq 2d}$. The following statements hold:
  1. (i)
    There exists a linear operator
    $A:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
    and a constant $\tau\geq 0$ such that
    $e^{tA}S\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
    for all $t\geq\tau$.
  2. (ii)
    Let
    $f\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
    and let
    $l:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R},\quad l(p):=\int_{K}p(x)\cdot e^{-x^{2}}\leavevmode\nobreak\ \mathrm{d}x.$
    Then
    $A:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d},\quad p\mapsto Ap:=l(p)\cdot f$
    is an operator as in (i).
  3. (iii)
    Let
    $L:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}$
    be a linear functional and let $A$ be an operator as in (i). Set
    $\tilde{S}:=e^{\tau A}S\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
    and
    $\tilde{L}:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R},\quad p\mapsto\tilde{L}(p):=L(e^{-\tau A}p).$
    Then
    $\inf_{p\in S}L(p)=\inf_{p\in\tilde{S}}\tilde{L}(p)\qquad\text{and}\qquad\sup_{p\in S}L(p)=\sup_{p\in\tilde{S}}\tilde{L}(p).$
Proof.
(i) + (ii): Since $C=\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$ is a full dimensional cone and $l(p)>0$ for all $p\in(S\cup C)\setminus\{0\}$, 3.2 gives the assertion.
(iii): Follows from the identity $e^{-\tau A}e^{\tau A}=\mathds{1}$. ∎
Remark 3.4.
An equivalent results as in 3.3 for the sums of squares cone also holds for the Waring cone by 3.2. $\circ$
Since 3.3 contains the degree bound $\leq 2d$, it can also be interpreted to work on the homogeneous polynomials:
$\mathds{R}[x_{0},x_{1},\dots,x_{n}]_{=2d}\quad\cong\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}.$
Note, the construction of the operator $A$ in 3.2 and 3.3 is very specific, since $A$ is made in such a way that $f\cdot[0,\infty)$ becomes an accumulation ray, i.e., with increasing time $t$ every $e^{tA}p$ converges towards the ray spanned by $f$. This requires the degree bounds in these results. In the following theorem the degree restriction is removed for $K=\mathds{R}^{n}$. It is our second main result.
Theorem 3.5.
Let $n\in\mathds{N}$. The following statements hold:
  1. (i)
    There exists a linear operator
    $A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
    with
    $A\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
    for all $d\in\mathds{N}_{0}$ such that
    $e^{A}\mathrm{Pos}(\mathds{R}^{n})_{\leq 2d}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
    for all $d\in\mathds{N}_{0}$ and therefore
    $e^{A}\mathrm{Pos}(\mathds{R}^{n})\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
  2. (ii)
    For $d\in\mathds{N}_{0}$, define the linear operators $l_{d}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}$ by
    $p=\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}x^{\alpha}\quad\mapsto\quad l_{d}(p):=\int_{\mathds{R}^{n}}e^{-x^{2}}\cdot\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|=2d}p_{\alpha}x^{\alpha}\leavevmode\nobreak\ \mathrm{d}x.$ (4)
    Then for any sequence $(f_{d})_{d\in\mathds{N}_{0}}\subseteq\mathds{R}[x_{1},\dots,x_{n}]$ of polynomials with
    $\deg f_{d}=2d\quad\text{and}\quad f_{d}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
    there exist constants $c_{0},c_{1},c_{2},{\dots}>0$ such that the operator $A$ in (i) is of the form
    $A=\sum_{d\in\mathds{N}_{0}}c_{d}\cdot l_{d}\cdot f_{d},\qquad\text{i.e.,}\qquad Ap=\sum_{d\in\mathds{N}_{0}}c_{d}\cdot l_{d}(p)\cdot f_{d}$
    for all $p\in\mathds{R}[x_{1},\dots,x_{n}]$.
Proof.
It is sufficient to prove (ii). We prove the statement via induction over the degree $d$.
$d=0$: Since
$\mathrm{Pos}(\mathds{R}^{n})_{\leq d=0}=\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d=0}^{2}=[0,\infty)\cdot 1,$
we have after rescaling $f_{0}=1$ and we can set $c_{0}=1$ to get
$A_{0}:=l_{0}\cdot 1.$
$d\to d+1$: For $d\in\mathds{N}_{0}$, assume we have an operator
$A_{d}=\sum_{k=0}^{d}c_{k}\cdot l_{k}\cdot f_{k}$
with
$A_{d}\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2k}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2k}$
and
$e^{A}\mathrm{Pos}(\mathds{R}^{n})_{\leq 2k}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq k}^{2}$
for all $k=0,1,\dots,d$. The operator
$A_{d}:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d+2}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d+2}$
is a bounded operator. Furthermore, $\mathrm{Pos}(\mathds{R}^{n})_{\leq 2d+2}$ is a pointed cone with compact base. For any
$p=\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}x^{\alpha}\in\mathrm{Pos}(\mathds{R}^{n})$
with $\deg p=2d+2$ the homogeneous part
$p|_{=2d+2}=\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|=2d+2}p_{\alpha}\cdot x^{\alpha}$
is non-negative, i.e., $p|_{=2d+2}\geq 0$ on $\mathds{R}^{n}$, and therefore $l_{d+1}(p)>0$. Hence, for
$A_{d+1}:=A_{d}+c_{d+1}\cdot l_{d+1}\cdot f_{d+1}=\sum_{k=0}^{d+1}c_{k}\cdot l_{k}\cdot f_{k}$
there exists a $c_{d+1}\gg 0$ such that
$e^{A_{d+1}}\mathrm{Pos}(\mathds{R}^{n})_{\leq 2d+2}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d+1}^{2},$
since
$f_{d+1}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d+1}^{2}$
and
$e^{A_{d}+c_{d+1}\cdot l_{d+1}\cdot f_{d+1}}:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d+2}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d+2}$
is continuous in $c_{d+1}$. That is the same fixed-ray argument as in 3.2 and $A_{d}$ is only a perturbation to $c_{d+1}\cdot l_{d+1}\cdot f_{d+1}$, since $c_{d+1}\gg c_{0},\dots,c_{d}$.
Now, set
$A:=\sum_{d\in\mathds{N}_{0}}c_{d}\cdot l_{d}\cdot f_{d}.$
Then $A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ is a well-defined linear operator, since $l_{D}(p)=0$ for all $p\in\mathds{R}[x_{1},\dots,x_{n}]$ and $D\in\mathds{N}_{0}$ such that $\deg p<2D$, i.e., the sum
$Ap=\sum_{d\in\mathds{N}_{0}}c_{d}\cdot l_{d}(p)\cdot f_{d}$
contains only finitely many non-zero terms. This proves (ii). ∎
Remark 3.6.
In the previous theorem we can, of course, rescale the given
$f_{d}=\sum_{\alpha\in\mathds{N}_{0}^{n}}f_{d,\alpha}\cdot x^{\alpha}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}.$
Hence, we can assume without loss of generality
$\sum_{\alpha\in\mathds{N}_{0}^{n}}f_{d,\alpha}^{2}=1.$
Then, in general,
$0\quad\ll\quad c_{0}\quad\ll\quad c_{1}\quad\ll\quad{\dots}\quad\ll\quad c_{d}\quad\ll\quad{\dots}\ .$
This increase in the constants $c_{d}$ reflects the fact that “there are significantly more non-negative polynomials than sums of squares” [3]. $\circ$
Remark 3.7.
From 3.5 we find that given any
$S\subseteq\mathrm{Pos}(\mathds{R}^{n})$
and any linear functional
$L:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},$
the optimization problems
$\inf_{p\in S}L(p)\qquad\text{and}\qquad\sup_{p\in S}L(p)$
transform into optimization problems
$\inf_{p\in S}L(p)\ =\ \inf_{p\in e^{A}S}\tilde{L}(p)\qquad\text{and}\qquad\sup_{p\in S}L(p)\ =\ \sup_{p\in e^{A}S}\tilde{L}(p)$
over the subset $e^{A}S$ of sums of squares $\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$ with
$\tilde{L}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},\quad p\mapsto\tilde{L}(p):=L(e^{-A}p).$
Here, $A\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\subseteq\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$ for all $d\in\mathds{N}_{0}$ implies
$e^{A}\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
for all $d\in\mathds{N}_{0}$, i.e., no additional degree blow up is introduced and also the Lasserre hierarchy can be used. $\circ$
A close look at the proof of 3.5 reveals that the operator $A$ in 3.5 (ii) can be constructed because the $l_{d}$ are given by (4). If similar linear functionals exist for $\mathrm{Pos}(K)$ for a closed $K\subseteq\mathds{R}^{n}$, then we get the following.
Corollary 3.8.
Let $n\in\mathds{N}$ and let $K\subseteq\mathds{R}^{n}$ be closed. Assume there exist
$0\quad\leq\quad d_{0}\quad<\quad d_{1}\quad<\quad d_{2}\quad<\quad{\dots}$
in $\mathds{N}_{0}$ and linear functionals $l_{i}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}$ with
$l_{i}(p)$ $>0$ for $\text{for all}\quad p\in\mathrm{Pos}(K)_{\leq 2d_{i}}\setminus\mathrm{Pos}(K)_{\leq 2d_{i-1}}$
and
$l_{i}(x^{\alpha})$ $=0$ for $\text{for all}\quad\alpha\in\mathds{N}_{0}^{n}\ \text{with}\ |\alpha|>2d_{i}$
for all $i\in\mathds{N}_{0}$. Then for any sequence $(f_{i})_{i\in\mathds{N}_{0}}$ of polynomials with
$\deg f_{i}=2d_{i}\quad\text{and}\quad f_{i}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d_{i}}^{2}$
there exist constants
$0\quad<\quad c_{0}\quad<\quad c_{1}\quad<\quad c_{2}\quad<\quad{\dots}$
such that the linear operator
$A=\sum_{i\in\mathds{N}_{0}}c_{i}\cdot l_{i}\cdot f_{i}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
is well-defined with
$e^{A}\mathrm{Pos}(K)_{\leq 2d_{i}}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d_{i}}^{2}$
for all $i\in\mathds{N}_{0}$ and therefore
$e^{A}\mathrm{Pos}(K)\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
Proof.
Proceed as in the proof of 3.5. ∎
In 3.5 and 3.8 the conditions
$\deg f_{d}=2d\qquad\text{and}\qquad\deg f_{i}=2d_{i}$
are already implied by
$f_{d}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}\qquad\text{and}\qquad f_{i}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d_{i}}^{2}.$
However, they are included for clarity.
The conditions in 3.8 are fulfilled if $K\subseteq\mathds{R}^{n}$ contains a full-dimensional cone.
Theorem 3.9.
Let $n\in\mathds{N}$, let $K\subseteq\mathds{R}^{n}$, and let $(f_{d})_{d\in\mathds{N}_{0}}$ be a sequence of polynomials with
$\deg f_{d}=2d\quad\text{and}\quad f_{d}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$
for all $d\in\mathds{N}$. If $K$ contains a full-dimensional cone $G$, then define
$l_{d}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},\quad\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}\cdot x^{\alpha}\mapsto\int_{G}e^{-x^{2}}\cdot\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|=2d}p_{\alpha}x_{\alpha}\leavevmode\nobreak\ \mathrm{d}x$
for all $d\in\mathds{N}_{0}$ and
$\tilde{l}_{d}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},\quad\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}\cdot x^{\alpha}\mapsto\int_{G}e^{-x^{2}}\cdot\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|=2d-1}p_{\alpha}x_{\alpha}\leavevmode\nobreak\ \mathrm{d}x,$
for all $d\in\mathds{N}$, and then there exist coefficients $c_{0},\tilde{c}_{1},c_{1},\tilde{c}_{2},c_{2},{\dots}\geq 0$ such that
$A:=\sum_{k=0}^{\infty}c_{k}\cdot l_{k}\cdot f_{k}+\sum_{k=1}^{\infty}\tilde{c}_{k}\cdot\tilde{l}_{k}\cdot f_{k}$
fulfills
$e^{A}\mathrm{Pos}(K)_{\leq 2d-1}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
for all $d\in\mathds{N}$, and
$e^{A}\mathrm{Pos}(K)_{\leq 2d}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
for all $d\in\mathds{N}_{0}$, i.e.,
$e^{A}\mathrm{Pos}(K)\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
Proof.
We proceed as in 3.5 by induction over the degree $d$. By taking only a full-dimensional subcone $B$ of $G$ spanned by the linearly independent vectors $b_{1},\dots,b_{n}$ and then by a linear variable transformation $x_{1},\dots,x_{n}$ to $\tilde{x}_{1},\dots,\tilde{x}_{n}$ such that $\tilde{b}_{1},\dots,\tilde{b}_{n}$ are the standard unit vectors of $\mathds{R}^{n}$, we can assume without loss of generality that $G=[0,\infty)^{n}$.
For $d=0$, let $A_{0}:=l_{0}\cdot 1$. Then
$e^{A_{0}}\mathrm{Pos}(K)_{\leq 0}=\mathrm{Pos}(K)_{\leq 0}=[0,\infty)\cdot 1=\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq 0}^{2}.$
Assume, for $d\in\mathds{N}_{0}$, we already have an operator
$A_{d}:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
with
$A_{d}=\sum_{k=0}^{d}c_{k}\cdot l_{k}\cdot f_{k}+\sum_{k=1}^{d}\tilde{c}_{k}\cdot\tilde{l}_{k}\cdot f_{k}$
such that
$e^{A_{d}}\mathrm{Pos}(K)_{\leq 2k-1}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq k}^{2}$
for all $k=1,\dots,d$, and
$e^{A_{d}}\mathrm{Pos}(K)_{\leq 2k}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq k}^{2}$
for all $k=0,\dots,d$. Let $p\in\mathrm{Pos}(K)$ with $\deg p=2d+1$, i.e.,
$p=\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|\leq 2d+1}p_{\alpha}x^{\alpha}.$
Then
$p|_{2d+1}=\sum_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|=2d+1}p_{\alpha}x^{\alpha}\geq 0$
on $K$. Let $f_{d+1}\in\mathrm{int}\,\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d+1}^{2}$ with
$\tilde{l}_{d+1}(f_{d+1})>0\quad\text{and}\quad l_{d+1}(f_{d+1})>0.$
Since $e^{A_{d}+\tilde{c}_{d+1}\tilde{l}_{d+1}f_{d+1}}$ is continuous in $\tilde{c}_{d+1}$ and since $\mathrm{Pos}(K)_{\leq 2d+1}$ is pointed and closed, i.e., it has a compact base, there exists a $\tilde{c}_{d+1}\geq 0$ such that
$e^{A_{d}+\tilde{c}_{d+1}\tilde{l}_{d+1}f_{d+1}}\mathrm{Pos}(K)_{\leq 2d+1}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d+1}^{2}.$
Since $e^{A_{d}+\tilde{c}_{d+1}\tilde{l}_{d+1}f_{d+1}+c_{d+1}l_{d+1}f_{d+1}}$ is continuous in $c_{d+1}$, $\mathrm{Pos}(K)_{\leq 2d+2}$ is closed and pointed, i.e., has a compact base, there exists a $c_{d+1}\geq 0$ such that
$\tilde{c}_{d+1}\tilde{l}_{d+1}(p)+c_{d+1}l_{d+1}(p)\geq 0$
for all $p\in\mathrm{Pos}(K)_{\leq 2d+2}$ and
$e^{A_{d}+(\tilde{c}_{d+1}\tilde{l}_{d+1}+c_{d+1}l_{d+1})\cdot f_{d+1}}\mathrm{Pos}(K)_{\leq 2d+2}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d+1}^{2}.$
At last, set
$A:=\sum_{k=0}^{\infty}c_{k}\cdot l_{k}\cdot f_{k}+\sum_{k=1}^{\infty}\tilde{c}_{k}\cdot\tilde{l}_{k}\cdot f_{k}.$
Then
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
is well-defined, since, for every $p\in\mathds{R}[x_{1},\dots,x_{n}]$, $\deg p<\infty$ and hence $Ap$ contains only finitely many non-zero terms. Since
$A\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
for all $d\in\mathds{N}_{0}$, $e^{A}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ is well-defined with
$e^{A}\mathrm{Pos}(K)_{\leq 2d-1}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
for all $d\in\mathds{N}$,
$e^{A}\mathrm{Pos}(K)_{\leq 2d}\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}^{2}$
for all $d\in\mathds{N}_{0}$, and hence
$e^{A}\mathrm{Pos}(K)\quad\subseteq\quad\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$
which ends the proof. ∎
If the assumptions in 3.8 are fulfilled, then an equivalent remark as 3.7 holds for $\mathrm{Pos}(K)$ instead of $\mathrm{Pos}(\mathds{R}^{n})$.
Remark 3.10.
If in 3.5 (or 3.8 or 3.9) each $f_{d}$ is in the interior of the Waring cone, i.e., the cone of even powers of linear forms of degree $2d$, then there exists a linear operator $A$ as in 3.5 (resp. 3.8) such that $e^{A}\mathrm{Pos}(\mathds{R}^{n})$ (resp. $e^{A}\mathrm{Pos}(K)$) is degree preserving and is a subset of the Waring cone. $\circ$
3.9 does not hold for compact $K\subseteq\mathds{R}^{n}$. The following 3.11 shows that, for compact $K\subseteq\mathds{R}^{n}$, there exists no linear operator
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
such that
$e^{A}\mathrm{Pos}(K)\ \subseteq\ \sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
3.11 is even stronger: There exists no bijective linear operator
$T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
with
$T\mathrm{Pos}(K)\ \subseteq\ \sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
Theorem 3.11.
Let $n\in\mathds{N}$ and let $K\subseteq\mathds{R}^{n}$ be compact. There exists no bijective and linear operator
$T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
such that
$T\mathrm{Pos}(K)\ \subseteq\ \sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
Proof.
Assume there exists a bijective linear operator
$T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]\quad\text{with}\quad T\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
Since $T$ is bijective, there exist $p\in\mathrm{Pos}(K)$ and $c>0$ with
$c-p\in\mathrm{Pos}(K)\qquad\text{and}\qquad\deg Tp\ >\ \deg T1.$
Then
$Tc=T(p+c-p)=Tp+T(c-p)$
with
$Tc,\ Tp,\ T(c-p)\quad\in\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$
implies
$\deg Tc\ =\ \max\{\deg Tp,\ \deg T(c-p)\}\ \geq\ \deg Tp\ >\ \deg Tc$
which is a contradiction. ∎
4 The set $\mathfrak{g}$ and its regular Fréchet Lie Algebras
In 3.2 and 3.3 we gave explicit operators
$A:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
such that certain given sets $S$ are moved by $e^{tA}$ to the nicer sets $e^{tA}S$, e.g. subsets of the sums of squares cone or the Waring cone. These results are restricted to the degrees $d$ and $2d$, i.e., for every $d\in\mathds{N}$ another operator $A$ and constant $\tau$ in 3.3 (iii) must be calculated. In 3.5 we gave an explicit linear operator $A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ with
$e^{A}\mathrm{Pos}(\mathds{R}^{n})\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}.$
For $\mathrm{Pos}(K)$ with closed $K\subseteq\mathds{R}^{n}$ we needed additional assumptions in 3.8. For compact $K\subseteq\mathds{R}^{n}$, in 3.11 we show that no operator $A$ with
$e^{A}\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$
exists. We therefore study now linear operators
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
such that
$e^{tA}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
is well-defined for all $t\in\mathds{R}$.
Definition 4.12.
Let $n\in\mathds{N}$. We define
$\mathfrak{g}:=\Big{\{}A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]\ \text{linear}\,\Big{|}\\ e^{tA}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]\ \text{well-defined for all}\ t\in\mathds{R}\Big{\}}.$
It is sufficient in the 4.12 of $\mathfrak{g}$ that $e^{tA}$ is well-defined as a linear map $\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$ only for all $t\in(-\varepsilon,\varepsilon)$ for some $\varepsilon>0$.
Clearly,
$\mathfrak{g}=\mathds{R}\cdot\mathfrak{g},$
i.e., $\mathfrak{g}$ is a cone and $\mathfrak{g}=-\mathfrak{g}$.
However, $\mathfrak{g}$ is not convex and it is not a vector space, i.e., there are $A,B\in\mathfrak{g}$ such that $A+B\notin\mathfrak{g}$. We have the following example of such maps $A$ and $B$.
Example 4.13.
Let $n=1$. Define the linear operators
$A:\mathds{R}[x]\to\mathds{R}[x]\qquad\text{and}\qquad B:\mathds{R}[x]\to\mathds{R}[x]$
by
$Ax^{k}:=\begin{cases}x^{k}&\text{for}\ k=2m\ \text{with}\ m\in\mathds{N}_{0},\\ x^{k+1}&\text{for}\ k=2m-1\ \text{with}\ m\in\mathds{N}\end{cases}$
and
$Bx^{k}:=\begin{cases}x^{k+1}&\text{for}\ k=2m\ \text{with}\ m\in\mathds{N}_{0},\\ x^{k}&\text{for}\ k=2m-1\ \text{with}\ m\in\mathds{N}\end{cases}$
for all $k\in\mathds{N}_{0}$ with linear extension to all $\mathds{R}[x]$. Then
$A\mathds{R}[x]_{\leq 2m}\ \subseteq\ \mathds{R}[x]_{\leq 2m}\qquad\text{and}\qquad B\mathds{R}[x]_{\leq 2m+1}\ \subseteq\ \mathds{R}[x]_{\leq 2m+1}$
for all $m\in\mathds{N}_{0}$. Since $A|_{\mathds{R}[x]_{\leq 2m}}$ and $B|_{\mathds{R}[x]_{\leq 2m+1}}$ are linear operators on finite dimensional spaces, i.e., matrices, $e^{tA}$ and $e^{tB}$ are well-defined for every $p\in\mathds{R}[x]$ and hence they are well-defined as linear maps $\mathds{R}[x]\to\mathds{R}[x]$: $A,B\in\mathfrak{g}$. However,
$(A+B)x^{k}\quad=\quad\begin{cases}x^{k}+x^{k+1}&\text{for}\ k=2m\ \text{with}\ m\in\mathds{N}_{0},\\ x^{k+1}+x^{k}&\text{for}\ k=2m-1\ \text{with}\ m\in\mathds{N}\end{cases}\quad=\quad x^{k}+x^{k+1}$
for all $k\in\mathds{N}_{0}$, i.e.,
$\deg\left(\sum_{i=0}^{d}\frac{t^{i}}{i!}\cdot(A+B)^{i}1\right)=d\quad\xrightarrow{d\to\infty}\quad\infty$
for all $t\neq 0$. Hence, $e^{t(A+B)}$ is not a map $\mathds{R}[x]\to\mathds{R}[x]$ and $A+B\notin\mathfrak{g}$. $\circ$
The previous example showed that if for a linear operator
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
there exist $d_{0}<d_{1}<d_{2}<\dots$ in $\mathds{N}_{0}$ with
$A\mathds{R}[x_{1},\dots,x_{n}]_{\leq d_{i}}\ \subseteq\ \mathds{R}[x_{1},\dots,x_{n}]_{\leq d_{i}}$ (5)
for all $i\in\mathds{N}_{0}$, then $A\in\mathfrak{g}$. This was the property used in [7, 9]. The next examples shows that (5) is sufficient, but (5) is not necessary.
Example 4.14.
Let $n=1$ and let $(p_{i})_{i\in\mathds{N}}=(2,3,5,7,11,\dots)$ be the list of prime numbers. Define the linear map $A:\mathds{R}[x]\to\mathds{R}[x]$ by
$Ax^{k}:=\begin{cases}x^{p_{m+1}}&\text{for}\ k=2m\ \text{with}\ m\in\mathds{N},\\ 0&\text{else}.\end{cases}$
Then
$A\mathds{R}[x]_{\leq d}\ \not\subseteq\ \mathds{R}[x]_{\leq d}$
for all $d\in\mathds{N}$ with $d\geq 8$. However,
$A1=0,\quad Ax=0,\quad Ax^{2}=x^{p_{2}}=x^{3},\quad Ax^{3}=0,\quad Ax^{4}=x^{p_{3}}=x^{5},\quad\dots$
and therefore
$A^{2}x^{k}=0$
for all $k\in\mathds{N}_{0}$, i.e., $e^{tA}:\mathds{R}[x]\to\mathds{R}[x]$ is well-defined for all $t\in\mathds{R}$ and $A\in\mathfrak{g}$. $\circ$
The following theorem gives characterizations of $\mathfrak{g}$.
Theorem 4.15.
Let $n\in\mathds{N}$ and let
$A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
be linear. The following are equivalent:
  1. (i)
    $A\in\mathfrak{g}$.
  2. (ii)
    $\sup_{k\in\mathds{N}_{0}}\deg A^{k}x^{\alpha}<\infty$ for all $\alpha\in\mathds{N}_{0}^{n}$.
  3. (iii)
    For all $i\in\mathds{N}_{0}$, there exist subspaces $V_{i}\subseteq\mathds{R}[x_{1},\dots,x_{n}]$ with
    1. (a)
      $\dim V_{i}<\infty$,
    2. (b)
      $\bigcup_{i\in\mathds{N}_{0}}V_{i}=\mathds{R}[x_{1},\dots,x_{n}]$, and
    3. (c)
      $AV_{i}\subseteq V_{i}$ for all $i\in\mathds{N}_{0}$.
Proof.
The implications “(ii) $\Rightarrow$ (i)” and “(iii) $\Rightarrow$ (i)” are clear.
(i) $\Rightarrow$ (ii): Since $e^{tA}$ is well-defined, for each $\alpha\in\mathds{N}_{0}^{n}$, there exists a degree $D=D(\alpha)\in\mathds{N}_{0}^{n}$ such that
$e^{tA}x^{\alpha}\ \subseteq\ \mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$
for all $t\in\mathds{R}$. Hence,
$A^{k}x^{\alpha}\ =\ \partial^{k}_{t}e^{tA}x^{\alpha}\Big{|}_{t=0}\ \in\ \mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$
for all $k\in\mathds{N}_{0}$, which proves (ii).
[(i) $\Leftrightarrow$ (ii)] $\Rightarrow$ (iii): Let $\alpha\in\mathds{N}_{0}^{n}$. By (i) and (ii), there exists a $D=D(\alpha)\in\mathds{N}_{0}$ with
$e^{tA}x^{\alpha},\ A^{k}x^{\alpha}\ \in\ \mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$
for all $k\in\mathds{N}_{0}$ and all $t\in\mathds{R}$. Set
$V_{\alpha,0}:=\mathds{R}\cdot x^{\alpha}.$
By (ii),
$AV_{\alpha,0}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq D}.$
For all $i\in\mathds{N}_{0}$, define
$V_{\alpha,i+1}:=V_{\alpha,i}+AV_{\alpha,i},$
i.e.,
$V_{\alpha,k}=\mathds{R}\cdot x^{\alpha}+\mathds{R}\cdot Ax^{\alpha}+\dots+\mathds{R}\cdot A^{k}x^{\alpha}$
for all $k\in\mathds{N}_{0}$. By (ii) and the definition of $V_{\alpha,i+1}$,
$V_{\alpha,i}\quad\subseteq\quad V_{\alpha,i+1}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$
and hence
$V_{\alpha,0}\quad\subseteq\quad V_{\alpha,1}\quad\subseteq\quad\dots\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq D}.$ (6)
Since (6) is an increasing sequence of finite dimensional vector spaces $V_{\alpha,i}$ bounded from above by the finite dimensional vector space $\mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$, there exists an index $I(\alpha)\in\mathds{N}_{0}$ such that
$V_{\alpha,I(\alpha)}=V_{\alpha,I(\alpha)+1},$
i.e.,
$\dim V_{\alpha,I(\alpha)}<\infty\qquad\text{and}\qquad AV_{\alpha,I(\alpha)}\subseteq V_{\alpha,I(\alpha)}.$
Since $\alpha\in\mathds{N}_{0}^{n}$ was arbitrary and $x^{\alpha}\in V_{\alpha,I(\alpha)}$,
$\mathds{R}[x_{1},\dots,x_{n}]=\bigcup_{\alpha\in\mathds{N}_{0}^{n}}V_{\alpha,I(\alpha)}.$
Since $\mathds{N}_{0}^{n}$ is countable, we proved (iii) (a) – (c). ∎
4.15 tells us that $e^{tA}$ is only the matrix exponential function, since for an operator $A\in\mathfrak{g}$ we only need to know $A$ on every finite dimensional invariant subspace $V_{i}$, i.e., its calculation is easy as soon as the $V_{i}$ for $A$ are known. An algorithm for the calculation of these $V_{i}$ is explicitly given in the proof of 4.15.
For the invariant subspaces $V_{i}$ of $A\in\mathfrak{g}$ we have the following result.
Corollary 4.16.
Let $n\in\mathds{N}$, let $A\in\mathfrak{g}$, and let $V\subseteq\mathds{R}[x_{1},\dots,x_{n}]$ be a finite or infinite dimensional subspace. The following are equivalent:
  1. (i)
    $AV\subseteq V$.
  2. (ii)
    $e^{tA}V\subseteq V$ for all $t\in\mathds{R}$.
Proof.
The direction “(i) $\Rightarrow$ (ii)” follows from $e^{tA}=\sum_{k\in\mathds{N}_{0}}k!^{-1}\cdot t^{k}\cdot A^{k}$ and the direction “(ii) $\Rightarrow$ (i)” follows from $A=\partial_{t}e^{tA}\big{|}_{t=0}$. ∎
We have seen in 4.13 that $\mathfrak{g}$ is not closed under addition. The following example shows that $\mathfrak{g}$ is also not closed under multiplication.
Example 4.17 (4.13 continued).
Let $A,B:\mathds{R}[x]\to\mathds{R}[x]$ be given as in 4.13. Then
$ABx^{2m}=x^{2m+2},\qquad ABx^{2m+1}=x^{2m+2},$ (7)
and
$BAx^{2m}=x^{2m+1},\qquad BAx^{2m+1}=x^{2m+3},$ (8)
for all $m\in\mathds{N}_{0}$, i.e., by 4.15 (ii), $AB\notin\mathfrak{g}$ as well as $BA\notin\mathfrak{g}$. $\circ$
If two operators $A,B\in\mathfrak{g}$ possess a common family $\{V_{i}\}_{i\in\mathds{N}_{0}}$ of invariant and finite dimensional subspace $V_{i}$ as provided by 4.15 (iii), then
$A+B\in\mathfrak{g},\quad AB\in\mathfrak{g},\quad\text{and}\quad BA\in\mathfrak{g}$
[…]
$[A,B]:=AB-BA\ \in\mathfrak{g}.$
In general, $\mathfrak{g}$ is not closed under this Lie bracket. The operators $A,B\in\mathfrak{g}$ in 4.13 provide an example of $[A,B]\notin\mathfrak{g}$.
Example 4.18 (4.13 and 4.17 continued).
Let $A:\mathds{R}[x]\to\mathds{R}[x]$ and $B:\mathds{R}[x]\to\mathds{R}[x]$ be defined as in 4.13. By (7) and (8) in 4.17,
$[A,B]x^{2m}=x^{2m+2}-x^{2m+1}\qquad\text{and}\qquad[A,B]x^{2m+1}=x^{2m+2}-x^{2m+3}$
for all $m\in\mathds{N}_{0}$, i.e., $[A,B]\notin\mathfrak{g}$ by 4.15 (ii). $\circ$
In summary, for $A,B\in\mathfrak{g}$, in order to have
$A+B\in\mathfrak{g},\quad AB\in\mathfrak{g},\quad BA\in\mathfrak{g},\quad\text{and}\quad[A,B]\in\mathfrak{g}$
it is sufficient that the operators $A$ and $B$ possess a common family $\{V_{i}\}_{i\in\mathds{N}_{0}}$ of invariant finite dimensional subspaces $V_{i}$ in 4.15 (iii).
Corollary 4.19.
Let $n\in\mathds{N}$ and let $\mathcal{V}:=\{V_{i}\}_{i\in\mathds{N}_{0}}$ be a family of subspaces $V_{i}\subseteq\mathds{R}[x_{1},\dots,x_{n}]$ such that
$\dim V_{i}<\infty\qquad\text{and}\qquad\mathds{R}[x_{1},\dots,x_{n}]=\bigcup_{i\in\mathds{N}_{0}}V_{i}.$
Then
$\mathfrak{g}_{\mathcal{V}}:=\big{\{}A\in\mathfrak{g}\,\big{|}\,AV_{i}\subseteq V_{i}\ \text{for all}\ i\in\mathds{N}_{0}\big{\}}$
is a regular Fréchet Lie algebra in $\mathfrak{g}$ with Lie bracket
$[\,\cdot\,,\,\cdot\,]:\mathfrak{g}_{\mathcal{V}}\times\mathfrak{g}_{\mathcal{V}}\to\mathfrak{g}_{\mathcal{V}},\quad(A,B)\mapsto[A,B]:=AB-BA.$
The set $G_{\mathcal{V}}:=\exp(\mathfrak{g}_{\mathcal{V}})$ is the corresponding regular Fréchet Lie group.
The question is, are all regular Fréchet Lie algebras in $\mathfrak{g}$ contained in some $\mathfrak{g}_{\mathcal{V}}$? For finite dimensional Lie algebras this is true.
Corollary 4.20.
Let $n\in\mathds{N}$ and let $\tilde{\mathfrak{g}}$ be a finite dimensional regular Fréchet Lie algebra in $\mathfrak{g}$. Then there exists a family $\mathcal{V}=\{V_{i}\}_{i\in\mathds{N}_{0}}$ of finite dimensional subspaces $V_{i}$ of $\mathds{R}[x_{1},\dots,x_{n}]$ with
$\mathds{R}[x_{1},\dots,x_{n}]=\bigcup_{i\in\mathds{N}_{0}}V_{i}$
such that $\tilde{\mathfrak{g}}\subseteq\mathfrak{g}_{\mathcal{V}}$.
Proof.
Let $g_{1},\dots,g_{N}$ be a vector space basis of $\tilde{\mathfrak{g}}$, i.e., $\dim\tilde{\mathfrak{g}}=N\in\mathds{N}$. Let
$a_{1},\dots,a_{N}\in\mathds{R}\quad\text{with}\quad a_{1}^{2}+\dots+a_{N}^{2}=1.$ (9)
Then for any $\alpha\in\mathds{N}_{0}^{n}$ there exists a $D=D(\alpha)\in\mathds{N}$ such that
$(a_{1}g_{1}+\dots+a_{N}g_{N})^{k}x^{\alpha}\quad\subseteq\quad\mathds{R}[x_{1},\dots,x_{n}]_{\leq D}$
for all $a_{1},\dots,a_{N}$ with (9), since $(a_{1}g_{1}+\dots+a_{N}g_{N})^{k}$ is continuous in $a_{1},\dots,a_{N}$ and the set (9) is compact. From here, continue as in the proof of 4.15 step “[(i) $\Leftrightarrow$ (ii)] $\Rightarrow$ (iii)”. ∎
An infinite dimensional regular Fréchet Lie algebra of $\mathfrak{g}$ does not need to be contained in some $\mathfrak{g}_{\mathcal{V}}$ as the following example shows.
Example 4.21.
Let $n\in\mathds{N}$, let $y\in\mathds{R}^{n}$, and let
$l_{y}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R},\qquad f\mapsto l_{y}(f):=f(y).$
Define
$\mathfrak{g}_{y}:=\big{\{}l_{y}\cdot p\,\big{|}\,p\in\mathds{R}[x_{1},\dots,x_{n}]\big{\}}.$
Then
  1. (i)
    $\mathfrak{g}_{y}\subseteq\mathfrak{g}$,
  2. (ii)
    $\alpha A+\beta B\in\mathfrak{g}_{y}$ for all $A,B\in\mathfrak{g}_{y}$ and $\alpha,\beta\in\mathds{R}$,
  3. (iii)
    $AB\in\mathfrak{g}_{y}$ for all $A,B\in\mathfrak{g}_{y}$,
  4. (iv)
    $[A,B]\in\mathfrak{g}_{y}$ for all $A,B\in\mathfrak{g}_{y}$, and
  5. (v)
    $\mathfrak{g}_{y}$ is closed,
i.e., $\mathfrak{g}_{y}$ is a regular Fréchet Lie algebra in $\mathfrak{g}$, since
$\alpha A+\beta B=l_{y}\cdot(\alpha p+\beta q)\qquad\text{and}\qquad AB=q(y)\cdot l_{y}\cdot p$
with $Af=f(y)\cdot p$, $Bf=f(y)\cdot q$, and $\alpha,\beta\in\mathds{R}$. To see that $\mathfrak{g}_{y}$ is closed let
$A_{k}=l_{y}\cdot p_{k}\qquad\text{with}\qquad A_{k}\xrightarrow{k\to\infty}\big{[}A:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]\big{]}.$
Hence, $A_{k}1=p_{k}\to p\in\mathds{R}[x_{1},\dots,x_{n}]$ as $k\to\infty$ and $A=l_{y}\cdot p\in\mathfrak{g}_{y}$. But since $\deg p$ in $A=l_{y}\cdot p$ is not bounded, $\mathfrak{g}_{y}$ is not contained in any $\mathfrak{g}_{\mathcal{V}}$. $\circ$
In 2.1 we gave the definition of a regular Fréchet Lie group and its Lie algebra. The Examples 4.13, 4.17, and 4.18 show that $\mathfrak{g}$ in the 4.12 is not a regular Fréchet Lie algebra of any regular Fréchet Lie group $G$. It is too large and not even a vector space. However, since it contains all operators $A$ such that $e^{tA}$ is well-defined, it contains all regular Fréchet Lie algebras in our case.
The set $\mathfrak{g}$ and its regular Fréchet Lie algebras are therefore not only of practical interest in the Theorems 3.2, 3.5, and 3.11, and 3.3, but they can also serve as simple toy examples to further study regular Fréchet Lie groups and their Lie algebras.
5 Three-Line-Proof of Stochel’s Theorem
Recall from Section 2, for a real sequence $s=(s_{\alpha})_{\alpha\in\mathds{N}_{0}^{n}}$ and a polynomial
$p(x)=\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}\cdot x^{\alpha}\quad\in\mathds{R}[x_{1},\dots,x_{n}],$
the Riesz functional $L_{s}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}$ is defined by
$L_{s}(p)=\sum_{\alpha\in\mathds{N}_{0}^{n}}p_{\alpha}\cdot s_{\alpha}.$
We get the following three-line-proof of Stochels Theorem [27].
Theorem 5.22.
Let $n\in\mathds{N}$ and $K\subseteq\mathds{R}^{n}$ closed. The following are equivalent:
  1. (i)
    $s$ is a $K$-moment sequence.
  2. (ii)
    For all $d\in\mathds{N}_{0}$, $s|_{\leq d}:=(s_{\alpha})_{\alpha\in\mathds{N}_{0}^{n}:|\alpha|\leq d}$ is a truncated $K$-moment sequence.
Proof.
(i) $\Rightarrow$ (ii): Clear.
[…]
Note, in (ii) actually a weaker property is used:
  1. (ii’)
    For all $d\in\mathds{N}_{0}$ and $p\in\mathrm{Pos}(K)_{\leq d}$, $L_{s|_{\leq}d}(p)\geq 0$.
The truncated sequences $s|_{\leq d}$ do not need to be truncated $K$-moment sequences.
Remark 5.23.
This three-line-proof was remarked by the author at the LAW25 conference. Jan Stochel was present in the lecture hall when the author made this remark. In the afternoon after this remark, Jan Stochel and Zenon Jablonski took the author aside and sat down on a bench in front of the Fakulteta za pomorstvo in promet Univerze v Ljubljani where the conference took place. The author, surrounded by Jan Stochel sitting on the right and Zenon Jablonski on the left, wrote down this three-line-proof. The original notes are shown in Figure 1.
Figure 1: The original handwritten notes by the author of the three-line-proof of Stochel’s Theorem [27], presented to Jan Stochel and Zenon Jablonski one afternoon at the LAW25 conference in front of the Fakulteta za pomorstvo in promet Univerze v Ljubljani.
While the author pointed out, that this three-line-proof must have been known before and that he was possibly not the first to notice that, Jan Stochel and Zenon Jablonski where not aware of this three-line-proof but immediately saw, that the weaker condition (ii’) is enough and that the theorem can be extended to graded algebras as long as an equivalent version of Haviland’s Theorem is known there, since the proof is based on the simple observations
$L_{s}(p)\overset{\deg p\leq d}{=}L_{s|{{}_{\leq}d}}(p)$
and
$\mathds{R}[x_{1},\dots,x_{n}]=\bigcup_{d\in\mathds{N}_{0}}\mathds{R}[x_{1},\dots,x_{n}]_{\leq d}.$
Both encouraged the author to publish this proof. Hence, it is included here. The original proof [27] also extends to *-semi-groups, where an analogue of Haviland’s Theorem is available, see e.g. [2].
Personally, I want to express my gratidute to both, Jan Stochel and Zenon Jablonski, for their kindness, openness, and impartiality to listen to me and to let me show this proof to them. While I was advised several times “not to fuck with the big guys” since they can hurt my career immensely, I am very happy to be able to meet Jan Stochel and Zenon Jablonski. It is refreshing that there are (still) mathematicians out there, who value good arguments and efficient proofs more than personal approval. I am very happy to personally meet them. $\circ$
6 Conclusions and Open Problem
In [7, 9, 8] we used the theory of regular Fréchet Lie groups and algebras to investigate linear maps
$T:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}]$
and
$e^{tA}:\mathds{R}[x_{1},\dots,x_{n}]\to\mathds{R}[x_{1},\dots,x_{n}].$
We extended previous results [11, 5, 16] about positivity preserver. Especially in [9] we claimed that the maps $e^{tA}$ could be used in linear optimization. This is shown in the current work by the Theorems 3.2 and 3.5.
In 3.3 we give for every $d\in\mathds{N}_{0}$ and subset $S\subseteq\mathrm{Pos}(K)_{\leq 2d}$ an explicit linear operator
$A:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$
such that $e^{tA}S$ is a subset of the sums of squares for all $t\geq\tau\geq 0$. In the optimization
$\min_{p\in S}L(p)$
over a linear functional $L:\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}\to\mathds{R}$, using the Schmüdgen or the Putinar Positivstellensatz can lead to degree blow ups in the $\sigma_{e}$ in (1) and (2), even though the original problem lives only in $\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d}$. This is circumvented by 3.3 (iii), since the Schmüdgen and the Putinar Positivstellensatz are removed.
But 3.2 and 3.3 depend on the fixed degree $d\in\mathds{N}_{0}$. In 3.5 we show that there exists an operator $A$ such that $e^{A}\mathrm{Pos}(\mathds{R}^{n})\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. But in 3.11 we see that, for compact $K\subseteq\mathds{R}^{n}$, there is not linear operator $A$ such that $e^{A}\mathrm{Pos}(A)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$. In fact, there is no bijective linear operator $T$ with $T\mathrm{Pos}(K)\subseteq\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$ for compact $K$.
An open problem of the current work is the following.
Open Problem 1.
Exists a (good or useful) description of the sets
$e^{\tau_{d_{i}}A}\big{(}S\cap\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d_{i}}\big{)}\qquad\text{or}\qquad e^{A}S$
in $\sum\mathds{R}[x_{1},\dots,x_{n}]_{\leq 2d_{i}}^{2}$ and $\sum\mathds{R}[x_{1},\dots,x_{n}]^{2}$, respectively?
In Section 4 we look at the structure of the set $\mathfrak{g}$ of all linear maps $A$ such that $e^{tA}$ is well-defined. We characterized all such maps $A$ in $\mathfrak{g}$. We showed with explicit examples that $\mathfrak{g}$ is not a vector space (4.13), $\mathfrak{g}$ is not closed under multiplication (4.17), and $\mathfrak{g}$ is not closed under the Lie bracket (4.18). We looked at special regular Fréchet Lie algebras in $\mathfrak{g}$ in 4.20 and 4.21. We have to work here in the full setting of non-commutative regular Fréchet Lie groups and algebras, where the theory is much more evolved than in the commutative case in [7].
With respect to 1 one also has to mention, that $\mathrm{Pos}(K)$ or general subsets $S\subseteq\mathrm{Pos}(K)$ are not spectrahedra or spectrahedral shadows [17]. This property is preserved under linear maps like $e^{tA}$, i.e., an NP-hard optimization problem remains NP-hard and “only” the degree blow ups in the Schmüdgen and the Putinar Positivstellensatz (1) and (2) are removed.
Acknowledgments
The author thanks Markus Schweighofer for introducing him to the concept of positivity preserver and especially giving him the reference [5]. The author thanks Jan Stochel and Zenon Jablonski for reading Section 5 and for the discussion at the LAW25 conference, where the author was invited by Aljaž Zalar to present the research [7, 9, 8] and the results in the present work.
Funding
The author is supported by the Deutsche Forschungsgemeinschaft DFG with the grant DI-2780/2-1 and his research fellowship at the Zukunftskolleg of the University of Konstanz, funded as part of the Excellence Strategy of the German Federal and State Government.
Conflict of Interest Statement
The author declares no conflict of interest.
Data Availability Statement
There is no associated data.